OpenStudy (anonymous):

Part 1: Simplify f(x) = 3x^2 - 4x +1/3x-1 Part 2: Using complete sentences, explain why f(1) = 0, f(0) = -1, and f(-1) = -2, yet f(1/3) is undefined. Show Work Please

OpenStudy (dan815):

factor the top

OpenStudy (dan815):

u get those valuse when you sub in those x values

OpenStudy (dan815):

the denominator has 3x-1 and the denominator cant be 0 so f(1/3) or when x=1/3 its un defined

OpenStudy (dan815):

3x-1=0 when x=1/3